OpenStudy (anonymous):

Can someone please tell me the integral of:

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2}y^2\sin(x)\]

OpenStudy (anonymous):

Well, not the answer, but the steps to get the answer

OpenStudy (freckles):

You have a mixture of variables there

OpenStudy (freckles):

we integrating w.r.t x or y?

OpenStudy (anonymous):

Sorry, this is multivariable calc so I only put in the part i needed help with. Wrt to x

OpenStudy (freckles):

If you are integrating w.r.t x you treat y like a constant

OpenStudy (anonymous):

\[\int\limits_{-3}^{3}\int\limits_{0}^{\pi/2}(y+y^2cosx)dxdy\]

OpenStudy (freckles):

do you know how to integrate sin(x) w.r.t x?

OpenStudy (anonymous):

Did I integrate \[y^2cosx\] properly?

OpenStudy (freckles):

\[\int\limits_{-3}^{3}[yx+y^2\sin(x)]|_{x=0}^{x=\frac{\pi}{2}} dy \]

OpenStudy (anonymous):

Okay, so I'm on the right track?

OpenStudy (freckles):

I don't know where you got integrating y^2*sin(x) w.r.t x from

OpenStudy (freckles):

Anyways you need to evaluate the limit part then integrate w.r.t y

OpenStudy (anonymous):

I probably didn't write it properly, but it looks like we're on the same path. Thank you.